OpenStudy (perii224):

PLEASE HELP WILL MEDAL The expression 2x(3x2 + 4) can be simplified into which of the following? 5x3 + 6x 6x2 + 8x 6x3 + 4 6x3 + 8x

OpenStudy (perii224):

@rishavraj Can you help me with this one too?

OpenStudy (perii224):

I think it's the second one

rishavraj (rishavraj):

how do u solve a(x + y) ??

rishavraj (rishavraj):

its the same :)) @perii224

OpenStudy (er.mohd.amir):

2x(3x^2+4) =2*x(3*x*x+4)=2*3*x*x*x+2*4*x=6x^3+8x

OpenStudy (perii224):

Ohh, so it's the 4th one?

rishavraj (rishavraj):

@perii224 yup

OpenStudy (perii224):

Thank youuu